# SequencePAR: Understanding Pedestrian Attributes via A Sequence Generation Paradigm

### v1

@@ IV-D Ablation Study
IV-E Visualization
In this section, we will give a visualization of the similarity matrix of learned pedestrian attributes. Then, we show the predicted pedestrian attributes using our proposed SequencePAR model.
@@ IV-E Visualization
Similarity Matrix.  As shown in Fig 3, we show the cosine similarity between various pedestrian attributes defined in the PETA datasets. We can find that some attributes are relatively independent, such as long hair, upper casual, upper formal, and upper jacket, as shown in the red rectangle. For attributes like age, shoes, attributes defined in the head region are highly correlated, as shown in blue rectangles. In our future works, we will consider learning the human attributes from the highly correlated groups.
Average Accuracy of Pedestrian Attribute.  As shown in Fig. 4, we compare the average accuracy of our method with VTB [18] on 20 pedestrian attributes on the PETA dataset. Our proposed method performs better than VTB [18] on most attributes, and some of the characteristics show significant improvement. Examples of these enhancements include “head sunglasses", “upper jackets", and “shoe sandals", which can improve by up to +12%. These results demonstrate the effectiveness of our generative model for pedestrian attribute recognition.
@@
IV-F Difference with Previous RNN-based PAR
Recurrent Neural Networks (RNN) are widely exploited in pedestrian attribute recognition for feature enhancement [55]. To be specific, Zhao et al. [55] propose to mine both the semantic and spatial correlations in attribute groups and predict the attributes in a group-by-group manner using the LSTM network. Wang et al. [15] propose the JRL model which can jointly learn the attribute context and correlation using the LSTM. Wang et al. [74] propose a recurrent memorized-attention module for multi-label image classification which is interpretable and contextualized. It contains a spatial transformer layer to locate attentional regions from the convolutional feature maps and an LSTM network to sequentially predict semantic labeling scores on the located regions. We can find that these works usually adopt RNN/LSTM for feature enhancement between different human parts, attentional regions mining, or correlations mining between human attributes.
In this paper, we propose a new Transformer-based generative prediction framework for pedestrian attribute recognition. The difference between our work and existing ones can be summarized as follows: 1). Existing works usually adopt the RNN/LSTM to model the attribute relations, while we adopt the advanced self-attention based Transformer network to achieve this goal. Therefore, our model performs better on parallelism, long-range dependencies, attention mechanisms, etc. 2). Many existing RNN/LSTM-based PAR models belong to the discriminative models, meanwhile, our proposed SequencePAR is a new generative framework for the pedestrian attribute recognition task. 3). Existing works predict the pedestrian attributes using the RNN/LSTM network which is trained from scratch, in contrast, our SequencePAR can make full use of the off-the-shelf pre-trained Transformer networks (e.g., vision model CLIP [24] and language model BERT [31]). Therefore, our proposed SequencePAR is significantly different from previous recurrent neural network based attribute recognition models. Also, according to the experimental results reported in sub-section IV-C, our model also exceeds the RNN based models on all the compared benchmark datasets.
IV-G Limitation Analysis
Although good performance can be achieved by our proposed SequencePAR model as validated in previous experiments, however, our model still performs poorly in some challenging scenarios. As illustrated in Fig. 6, some of our predicted pedestrian attributes are incorrect which are highlighted in red. It is easy to find that these images usually contain multiple persons or interfering objects which are really hard for the model to discriminate. Therefore, it will be a possible solution to first locate the true pedestrian we need to analyze, and then, predict the pedestrian attributes. We leave this as our future works.
@@ IV-G Limitation Analysis
V Conclusion
@@
In this work, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. It is proposed based on the fact that existing pedestrian attribute recognition (PAR) algorithms are following multi-label or multi-task learning. That is to say, their performance heavily depends on the specific classification heads and is easily influenced by imbalanced data or noisy samples. Therefore, we propose the SequencePAR which belongs to the generative models. To be specific, it extracts the pedestrian features using a pre-trained CLIP model and embeds the attribute set into query tokens under the guidance of text prompts. Then, a Transformer decoder is proposed to generate the human attributes by incorporating the visual features and attribute query tokens. The masked multi-head attention layer is introduced into the decoder module to prevent the model from remembering the next attribute while making attribute predictions during training. Extensive experiments on multiple widely used PAR datasets fully validated the effectiveness of our proposed SequencePAR. In our future work, we will consider introducing new main target pedestrian localization modules to eliminate the interference of background pedestrians, as discussed in Section IV-G.

### v2

@@ 4.4 Ablation Study
Figure 4: Visualization of Different Designs of The Masking Strategy. Gray represents the invisible.
Effectiveness of Causal Masks. As shown in Table 6, we first analyze the necessity of causal masking to determine the most effective masking strategy. We observe that removing the attention mask from the autoregressive decoder causes model failure. This is because we use the true values as input to reduce cumulative error during training, enabling the model to complete the entire prediction in one forward pass. Without the attention mask, the model can anticipate predictions through the attention mechanism, causing training to collapse. We then experiment with several masking strategies, including the ”group masking” strategy, where groups of attributes (e.g., head, top dress, bottom dress, etc.) are visible to each other, and causal masks are applied between attribute groups to model relationships. Sparse masking means the model cannot see the next K tokens. Experiments show that the optimal performance is achieved using the causal masking strategy. Specifically, the group masking strategy underperforms because elements within attribute groups can see future tokens, causing overfitting and significant performance degradation. Additionally, the sparse masking strategy generally performs poorly, especially with smaller mask step settings (e.g., Sparse Mask-1). Smaller mask lengths do not effectively obscure future tokens, leading to information leakage and causing the model to depend on future information, which negatively impacts training. However, increasing the mask length (e.g., Sparse Mask-10) improves performance. This is because, for attributes appearing later in the sequence, a larger mask length effectively obscures future tokens, leading to improved predictions. In summary, our results underscore the importance of the causal masking strategy. Causal masking prevents future information leakage, enhancing the model’s generalization ability.
Table 6: Comparison of Different Designs of The Masking Strategy on PETA dataset.
Masking Strategy PETA
Accuracy Precision Recall F1
Causal Mask 84.92 90.44 90.73 90.46
Group Mask 64.09 88.51 68.13 76.71
Sparse Mask-1 8.72 37.71 9.98 15.67
Sparse Mask-10 17.74 72.53 18.05 27.73
Sparse Mask-15 22.07 84.91 22.42 34.29
w/o Mask 7.18 37.33 7.52 12.28
Figure 5: Robustness Analysis under Label Noise and Long-tailed Distribution. (a) Performance degradation under 10% random label noise. (b) Performance after reducing tail attribute samples. (c) Attribute frequency distribution after reducing tail sampling.
Table 7: Comparison of Different Prompts for Attributes Expanding on PETA dataset.
Prompts PETA
Accuracy Precision Recall F1
<<CLASS>> 84.23 89.88 90.32 89.98
A photo of a <<CLASS>> 84.53 90.15 90.49 90.19
Custom Template 84.92 90.44 90.73 90.46
Table 8: Comparison of Different orders for Expanding on the PETA dataset.
Prompts PETA
Accuracy Precision Recall F1
Init 84.92 90.44 90.73 90.46
Inverse 85.01 90.41 90.89 90.53
Shuffle 1 84.69 90.23 90.49 90.25
Shuffle 1 84.62 90.25 90.42 90.21
Effectiveness of Prompts for Attributes Expanding. As shown in Table 7, we explored the effect of using two additional cues on the model’s performance. The template we used was more detailed and contextually appropriate, e.g., ”the age of this pedestrian is less than 40 years old”. Additionally, we explored two alternative cues. The first approach involved using the attribute words themselves directly. The results indicate a slight decrease in model performance across the four evaluation metrics (0.69/0.56/0.41/0.48), demonstrating that the cue template enhances attribute differentiation and mitigates the expression ambiguity of attribute words. The second approach involved using the commonly adopted CLIP template, ”A photo of a <<CLASS>>. The results show a slight decline in performance (0.39/0.29/0.24/0.27 for Acc/Prec/Recall/F1, respectively), further confirming that an appropriate semantic template aids the model in better understanding attribute semantics.
Table 9: Comparison of Model Performance under Original, Noisy, and Long-tailed Training Settings.
Train Set Methods PETA
Accuracy Precision Recall F1
Origin Ours 84.92 90.44 90.73 90.46
VTB[2] 79.60 86.76 85.69 86.71
PARFormer-L[8] 82.86 88.06 91.98 89.06
[…]
Analysing the Impact of Different Attribute Orders. To analyze the sequence relationships between attributes learned by the model, we randomly disrupted the PETA attribute list twice and compared the results with the original order. As shown in Table 8, the model still effectively captures the relationships between attributes with minimal deviation, even when their order is altered. However, disrupting the order led to a slight decrease in performance, likely due to the causal relationships between certain attributes. Across all metrics, performance dropped by approximately 0.2%. These results indicate that while the model demonstrates strong robustness to changes in attribute order, alterations to certain causal relationships still impact performance.
@@
Robustness Analysis under Label Noise and Long-tailed Distribution. As shown in Figure 5. To evaluate the robustness of our model against noisy annotations and imbalanced data, we conduct experiments involving random noise injection and long-tail sampling on the training dataset.
To assess the model’s ability to handle label noise, we introduce 10% label corruption into the training set. Specifically, each ground-truth attribute label was flipped with a probability of 10%, resulting in a noisy training set denoted as $D_{noise}$. We then retrain our model, the visual-text fusion method VTB, and the prior state-of-the-art method PARFormer on $D_{noise}$. As shown in Table 9. Under this noise setting, our model shows smaller performance drops across key metrics (Accuracy/Precision/Recall/F1: 9.26/8.84/2.18/5.39) compared to VTB (11.54/11.69/1.54/10.85). Although VTB exhibits a slightly smaller drop in Recall, our model maintains more balanced and stable Precision and Recall, resulting in superior overall F1 performance. Furthermore, compared to PARFormer [8], our model experiences smaller drops (5.16/3.78/3.30/2.18), highlighting the enhanced robustness provided by the generative framework. We attribute this to the autoregressive decoder, which models inter-attribute dependencies by conditioning each prediction on previously generated tokens, thus promoting context-aware denoising.
To further investigate robustness under extreme class imbalance, we construct a more severely long-tailed training set through attribute-level sampling. For each attribute, we identify tail classes with relatively few positive samples and randomly remove a fixed proportion of their positive labels. This increases the dataset’s mean imbalance ratio (Mean IR) [4], a commonly used metric in long-tailed classification, where a higher value indicates greater imbalance. Through this sampling strategy, we increase the Mean IR from 12.01 to 26.42 as illustrated in Figure 5, yielding a new long-tailed dataset denoted as $D_{lt}$.
All models exhibit performance degradation when trained on $D_{lt}$, however, our method demonstrates the smallest performance drop. As shown in Table 9, in terms of F1 score, SequencePAR drops only 2.78, while VTB and PARFormer-L drop 3.91 and 3.24, respectively. These results confirm the robustness and generalization capability of our method under both noisy and imbalanced training conditions.
4.5 Visualization
In this section, we provide a visualization of the similarity matrix for the learned pedestrian attributes. Additionally, we present the predicted pedestrian attributes generated by our proposed SequencePAR model.
Figure 7: Visualization of the attentions between attribute words during our model autoregression.
Similarity Matrix.  As shown in Fig 6, we show the cosine similarity between various pedestrian attributes defined in the PETA datasets. We can find that some attributes are relatively independent, such as long hair, upper casual, upper formal, and upper jacket, as shown in the red rectangle. For attributes like age and shoes, attributes defined in the head region are highly correlated, as shown in blue rectangles. In our future work, we will consider learning the human attributes from the highly correlated groups.
Attention Map Between Attribute Words.  We also present a visualization of the model’s attention during autoregression. As shown in Figure 7(a), the red box highlights the model’s increased attention to the Upper Formal attribute when predicting the Lower Formal attribute. Indeed, a causal relationship exists between these two attributes, and the model successfully captured this connection. Additionally, as shown in the yellow area of Figure 7(a), the model attends to both upper and lower body wear when predicting the Shoes Leather attribute, further demonstrating the effective utilization of contextual information by the model.
@@ 4.5 Visualization
Average Accuracy of Pedestrian Attribute.  As shown in Figure 8, we compare the average accuracy of our method with VTB [2] on 20 pedestrian attributes on the PETA dataset. Our proposed method performs better than VTB [2] on most attributes, and some of the characteristics show significant improvement. Examples of these enhancements include “head sunglasses”, “upper jackets”, and “shoe sandals”, which can improve by up to +12%. These results demonstrate the effectiveness of our generative model for pedestrian attribute recognition.
@@
Incorrectly Results.  Although our proposed SequencePAR model demonstrates strong performance, as validated in previous experiments, it still faces challenges in certain scenarios. As illustrated in Figure 10, some predicted pedestrian attributes are incorrect, highlighted in red. These mispredictions often occur in images containing multiple individuals or interfering objects, which are difficult for the model to distinguish. To address this, one potential solution is to first locate the target pedestrian before predicting their attributes. We present the prediction results for noisy samples from the PETA dataset. As shown in Figures 11 (a) and (c), the dataset incorrectly labels the pedestrian’s gender as male, resulting in a conflicting attribute pair of Long Hair and Male, which our model successfully corrects. Additionally, the dataset contains other types of noise, such as the contradictory attribute pair Head Muffler and Lower Shorts in Figure 11 (b), which violates seasonal consistency. As observed, our model mitigates the impact of such noisy samples by leveraging contextual information.
Figure 11: Visualization of some incorrectly predicted pedestrian attributes (The counterfactual attributes in the ground truth are represented in red, and the corrected predictions from our model are represented in bold.) using our proposed SequencePAR.
4.6 Difference with Existing Methods
We will discuss the differences between our method and existing methods. We categorize related methods into three types: PAR methods [2], sequential methods [36], and generative methods [3].
First, current mainstream multi-label classification-based [10, 20] PAR methods primarily treat pedestrian attribute prediction as multiple independent binary classification problems. These methods extract features using a backbone and input them into multiple classifiers to predict the confidence of each attribute. The PAR dataset contains a large number of negative samples, which often leads to low prediction confidence from the classifiers. Compared to the discriminative architectures used in mainstream PAR methods, our approach leverages a generative architecture to transform the independent prediction problem into a joint probability problem, thereby better utilizing the contextual information of attributes to logically predict errors (as shown in Figure 11). Additionally, our approach does not rely on the binary classifiers commonly used in PAR but instead predicts the index of attribute words in the vocabulary, alleviating the issue mentioned by Jia et al. [11], where the prediction confidence lies on the decision boundary as the number of attributes increases.
Second, RNN-based sequential methods [31] typically employ RNNs or LSTMs to sequentially model the dependencies among pedestrian attributes. However, these approaches still rely on discriminative prediction schemes, making them vulnerable to class imbalance issues. Our method transforms discriminative prediction into the task of predicting the index of pedestrian attributes, utilizing a joint positive sample-weighted NLL loss, which mitigates the issue of class imbalance. Moreover, our approach utilizes an advanced self-attention-based Transformer [29] network to model attribute relationships. It excels in parallelism, long-range dependencies, and attention mechanisms, with a more detailed analysis presented in Section 4.7.
Finally, compared with other generation-based methods such as image captioning [3], which typically employ a Transformer encoder-decoder architecture to perform autoregressive target description via masked Multi-Head Self-Attention (MHSA), our approach introduces several key innovations. Although MHSA has been widely used in tasks like image captioning, it has not yet been applied in pedestrian attribute recognition (PAR) to model the causal dependencies among attributes. Rather than directly adopting MHSA, we design a novel framework comprising an attribute query encoder, a sequence generation decoder, and a closed-vocabulary generation strategy optimized for structured attribute prediction. Experimental comparisons demonstrate that, when using the same ViT-L/14 visual encoder, our method outperforms discriminative models such as VTB* by more than 5% in accuracy on datasets like PETA. In addition, image captioning typically relies on an open vocabulary, which often requires substantial computational resources and diverse sentence-level annotations. Our model proposes an innovative closed attribute vocabulary, thereby reducing the likelihood of generating irrelevant tokens, accelerating convergence, and eliminating the need for additional sentence-level annotations.
4.7 Difference with Previous RNN-based PAR
Recurrent Neural Networks (RNN) are widely exploited in pedestrian attribute recognition for feature enhancement [36]. To be specific, Zhao et al. [36] propose to mine both the semantic and spatial correlations in attribute groups and predict the attributes in a group-by-group manner using the LSTM network. Wang et al. [31] propose a recurrent memorized-attention module for multi-label image classification which is interpretable and contextualized. The model includes a spatial Transformer layer to identify attentional regions from convolutional feature maps and an LSTM network to sequentially predict semantic labeling scores for these regions. These approaches commonly utilize RNNs//LSTMs for enhancing features across different human parts, mining attentional regions, or capturing correlations between human attributes.
In this paper, we propose a new Transformer-based generative prediction framework for pedestrian attribute recognition. The difference between our work and existing ones can be summarized as follows: 1). Existing works usually use the RNN//LSTM to model the attribute relations, while we adopt the advanced self-attention-based Transformer network to achieve this goal. Therefore, our model performs better on parallelism, long-range dependencies, attention mechanisms, etc. 2). Many existing RNN//LSTM-based PAR models belong to the discriminative models, meanwhile, our proposed SequencePAR is a new generative framework for the pedestrian attribute recognition task. 3). Existing works predict the pedestrian attributes using the RNN/ LSTM network which is trained from scratch. In contrast, our SequencePAR can make full use of the off-the-shelf pre-trained Transformer networks (e.g., the vision model CLIP [24] and language model BERT [13]). Therefore, our proposed SequencePAR is significantly different from previous recurrent neural network-based attribute recognition models. Also, according to the experimental results reported in sub-section 4.3, our model also exceeds the RNN-based models on all the compared benchmark datasets.
5 Conclusion and Future Works
@@
In this work, we propose a novel sequence generation paradigm for pedestrian attribute recognition, termed SequencePAR. It is proposed based on the fact that existing pedestrian attribute recognition (PAR) algorithms are following multi-label or multi-task learning. That is to say, their performance heavily depends on the specific classification heads and is easily influenced by imbalanced data or noisy samples. Therefore, we propose the SequencePAR which belongs to the generative models. To be specific, it extracts the pedestrian features using a pre-trained CLIP model and embeds the attribute set into query tokens under the guidance of text prompts. Then, a Transformer decoder is proposed to generate the human attributes by incorporating the visual features and attribute query tokens. The masked multi-head attention layer is introduced into the decoder module to prevent the model from remembering the next attribute while making attribute predictions during training. Extensive experiments on multiple widely used PAR datasets fully validated the effectiveness of our proposed SequencePAR.
This study reveals an unexpected phenomenon through systematic experimental analysis: in the pedestrian attribute prediction task under the generative paradigm, the performance of the beam search strategy is not significantly superior to that of the greedy search strategy. This phenomenon may arise from our closed-set attribute vocabulary design, which lacks synonyms, thus eliminating the need for the model to explore multiple candidate sequences during decoding. In our future work, we aim to advance the paradigm in two key dimensions. First, we will construct an open attribute vocabulary and enhance the model’s zero-shot prediction capabilities by integrating knowledge distillation techniques from multimodal large language models (MLLMs). This approach will facilitate generalized reasoning for previously unseen attributes. Second, we will employ Chain-of-Thought (CoT) reasoning techniques to improve the model’s interpretability in fine-grained attribute association reasoning and counterfactual causal inference while maintaining robust generalization to open-domain scenarios.